[SADT3] For scalar functions u and v, show that



B=( nabla u)*( nabla v)



is solenoidal and that



A= 1 2 (u nabla v-v nabla u)



is a vector potential for B, i.e. B= nabla* A

1
Expert's answer
2021-12-06T13:28:41-0500

"B=( \nabla u)\times( \nabla v)"

"A= \frac{1}{2} (u \nabla v-v \nabla u)"

"B= \nabla\times A"


for solenoidal vector field:

"\nabla \cdot B=0"


"\nabla u=\frac{\partial u}{\partial x}i+\frac{\partial u}{\partial y}j+\frac{\partial u}{\partial z}k"


"\nabla v=\frac{\partial v}{\partial x}i+\frac{\partial v}{\partial y}j+\frac{\partial v}{\partial z}k"


"B=( \nabla u)\times( \nabla v)=\begin{vmatrix}
 i & j&k \\
 \frac{\partial u}{\partial x}&\frac{\partial u}{\partial y}&\frac{\partial u}{\partial z}\\
\frac{\partial v}{\partial x}&\frac{\partial v}{\partial y}&\frac{\partial v}{\partial z}\\
\end{vmatrix}="


"=(\frac{\partial u}{\partial y}\frac{\partial v}{\partial z}-\frac{\partial v}{\partial y}\frac{\partial u}{\partial z})i-(\frac{\partial u}{\partial x}\frac{\partial v}{\partial z}-\frac{\partial u}{\partial z}\frac{\partial v}{\partial x})j+(\frac{\partial u}{\partial x}\frac{\partial v}{\partial y}-\frac{\partial u}{\partial y}\frac{\partial v}{\partial x})k=0"


so, "\nabla \cdot B=0"



"A= \frac{1}{2} (u \nabla v-v \nabla u)=\frac{1}{2} (u\frac{\partial v}{\partial x}i+u\frac{\partial v}{\partial y}j+u\frac{\partial v}{\partial z}k-v\frac{\partial u}{\partial x}i-v\frac{\partial u}{\partial y}j-v\frac{\partial u}{\partial z}k)"


"\nabla\times A=[\frac{\partial}{\partial y}(u\frac{\partial v}{\partial z}-v\frac{\partial u}{\partial z})-\frac{\partial}{\partial z}(u\frac{\partial v}{\partial y}-v\frac{\partial u}{\partial y})]i+[\frac{\partial}{\partial z}(u\frac{\partial v}{\partial z}-v\frac{\partial u}{\partial z})-\frac{\partial}{\partial x}(u\frac{\partial v}{\partial z}-v\frac{\partial u}{\partial z})]j+"


"+[\frac{\partial}{\partial x}(u\frac{\partial v}{\partial y}-v\frac{\partial u}{\partial y})-\frac{\partial}{\partial y}(u\frac{\partial v}{\partial x}-v\frac{\partial u}{\partial x})]k=B"